### v1

# MAMBO: High-resolution Generative Approach for Mammography Images

Abstract
Mammography is the gold standard for the detection and diagnosis of breast cancer. This procedure can be significantly enhanced with Artificial Intelligence (AI)-based software, which assists radiologists in identifying abnormalities. However, training AI systems requires large and diverse datasets, which are often difficult to obtain due to privacy and ethical constraints. To address this issue, the paper introduces MAMmography ensemBle mOdel (MAMBO), a novel patch-based diffusion approach designed to generate full-resolution mammograms. Diffusion models have shown breakthrough results in realistic image generation, yet few studies have focused on mammograms, and none have successfully generated high-resolution outputs required to capture fine-grained features of small lesions. To achieve this, MAMBO integrates separate diffusion models to capture both local and global (image-level) contexts. The contextual information is then fed into the final patch-based model, significantly aiding the noise removal process. This thoughtful design enables MAMBO to generate highly realistic mammograms of up to 3840×3840 pixels. Importantly, this approach can be used to enhance the training of classification models and extended to anomaly detection. Experiments, both numerical and radiologist validation, assess MAMBO’s capabilities in image generation, super-resolution, and anomaly detection, highlighting its potential to enhance mammography analysis for more accurate diagnoses and earlier lesion detection.
1 Introduction
Figure 1: Synthetic $38403840x$ 3840 mammogram generated using MAMBO. Best viewed when zoomed in.
Breast cancer is the most common malignancy and leading cause of cancer deaths in women around the world [6]. Mammography is the preferred method for the detection and prevention of breast cancer, as it is effective, cost-effective, and reliable, making it ideal for screening large populations. This, in turn, produces numerous mammograms, all of which should be analyzed by qualified medical professionals. Analysis often requires considerable effort, such as the need to detect subtle contrast variations caused by differences in breast density [37]. To address this challenge, radiologists have increasingly relied on computer-aided diagnosis (CAD) systems, which have become essential tools in mammography in screening centers and hospitals [4]. Advancements in Deep Learning (DL) have brought further improvements, setting a new standard in medical imaging, and enhancing both diagnosis accuracy and patient care [15]. Provided that sufficient diverse training data is available, a wide range of tasks can be tackled with DL [5]: image generation [22], risk prediction [42], cancer presence prediction [31], detection [24], segmentation [29], and anomaly detection [38].
However, high-quality medical images are difficult to obtain, often due to privacy and legal concerns. Thus, there is growing interest in DL-driven approaches for generating synthetic data [43], which can be thought of as a form of advanced data augmentation. One such approach is found in denoising diffusion models [10] which have recently been developed in radiology [20]. However, they typically support relatively low image resolutions (e.g., $256\times 256$ pixels), limiting their usefulness in mammography applications.
Mammography images are the highest resolution images used in radiology [26], ranging from $2300\times 1800$ pixels to $4096\times 5625$ pixels, where the size of a pixel corresponds to an area of 100 micron and 54 micron, respectively. This high resolution is needed to be able to observe cancers less than 1 cm in size, which has a significant effect on the outcome of the therapy. The 10-year survival rate drops from more 95% in patients in whom the observed cancer is less than 1 cm to about 60% when the cancer size is more than 3 cm [45].
To overcome these challenges, this paper introduces MAMmography ensemBle mOdel (MAMBO), a novel approach that leverages an ensemble of diffusion models to generate mammography images at their native resolution. MAMBO uses a patch-based approach for generation which is conditioned both on local and global (full-image) context. In a nutshell, MAMBO involves the development of three distinct models: the first generates “standard”-resolution ($256\times 256$ pixels) images to provide global context; the second increases the resolution to create local context for the target patch; and the third combines the outputs of both models to guide the generation of high-resolution patches, which are then reconstructed into a full-resolution synthetic mammogram. To the best of our knowledge, this is the first work to propose a diffusion model approach for generating very high-resolution synthetic mammograms.
The proposed approach is evaluated on three benchmarks: VinDr [17], RSNA [18] and InBreast [16]. This evaluation demonstrates the effectiveness of MAMBO in generating synthetic mammograms, considering common metrics adopted for generative models (e.g., Fréchet Inception Distance - FID), comparing the approach with previous methods, and expert radiologist evaluation. Example of a mammogram generated with MAMBO is shown in Fig. 1. Next, results for an anomaly detection downstream task are presented. This is achieved by generating a healthy counterpart for the input mammogram and comparing it to the original image. The approach makes pixel-wise predictions on the presence of a lesion without training on pixel-wise labels. Considering the scarcity of labeled medical data, an unsupervised approach is very valuable. Results as high as 0.423 are achieved in Intersection over union (IoU).
2 Related Work
Generating images in the medical domain. When it comes to generating mammography images, whether to train other models or as a foundation for anomaly detection, Generative Adversarial Networks (GANs) remain the core technology. Oyelade et al. [19] provided a comprehensive overview of the different approaches employed in this domain. The paper also proposed an approach for synthesizing specific regions of interest in mammography. While the approach can be considered patch-based, it can not generate whole mammograms. Park et al. [21] focused on generating whole images of healthy breasts and showed reasonable performance in terms of detecting cancer as anomalies [30]. This type of approach, similarly to the one proposed in this study, provides the opportunity to create pixel-level segmentation masks of lesions, using only weak labels in the form of image-level annotations. Using the StyleGAN2 architecture [12] for generation, an impressive FID score of 4.383 was achieved.
Recently, diffusion models have been widely adopted in computer vision [2], demonstrating impressive generative capabilities. As a result, there is growing research interest in applying these models to the field of radiology. For example, it has been shown to be effective for weakly-supervised anomaly detection [38], particularly when modifications are made to the diffusion process itself [40]. A study by Pan et al. showed that diffusion models are able to generate X-ray, MRI and CT images that are indistinguishable from the originals even for the eye of the professional radiologist [20] and therefore pass the visual Turing test. The study, however, did not use any mammography data, and the resulting images are of relatively low resolution ($256\times 256$ pixels).
When it comes to generative AI in medicine, one should mention the efforts directed towards a standardized baseline framework for it [22]. The models available, however, are not yet capable of generating high-resolution mammograms.
Generating high-resolution images. For high-resolution images, using the Latent Diffusion Model (LDM) is common [27]. This approach first compresses the image, typically using a trained Variational Autoencoder (VAE) [13] or a Vector-Quantized Variational Autoencoder (VQ-VAE) [34], thus transferring the image to a latent representation. Noising and denoising are applied to this latent representation, which is then transformed back into the pixel space. This approach helps to reduce the computational cost, which would be too high if the diffusion model were trained directly on high-resolution images. Unfortunately, training an LDM on mammography data requires large-scale datasets, and their lack of availability has hindered its development so far.
An alternative to LDM for high-resolution images is the patch-based approach, which is used in this study. While we are unaware of any applications of this method in mammography, Ding et al. recently proposed a technique called Patch-DM [3] for natural images. In the study, a sliding window approach is used with a local context provided by padding the patch and passing the overlap of the padding with the neighboring patches as context. It also proposes leveraging features extracted from a pretrained CLIP model [25] as global conditions, along with classifier-free guidance [9], to enhance training speed and stability.
Figure 2: In the three-stage approach of MAMBO, the first stage of the model generates a novel global context $\mathbf{x}_{0_{G}}$, which is then used to generate a set of local contexts $\mathbf{X}_{0_{L}}$ in the second stage, conditioned on shifted global contexts $\overrightarrow{\mathbf{X}_{0_{G}}}$. Synthetic global context and synthetic local contexts become the conditioning in the third stage to generate highly detailed patches $\mathbf{X}_{0_{P}}$, which are finally combined to obtain a high-resolution full mammogram.
3 Method
In this work, a novel approach is proposed to generate high-resolution synthetic mammography images using an ensemble of three diffusion models in a coarse to fine manner. This allows the number of parameters in each model to be kept roughly constant and at a reasonable amount for consumer-level hardware, while creating full-resolution results. Once all the patches of a full-resolution mammogram are generated, they are overlapped to generate output images without visible artifacts. The pipeline consists of three Denoising Diffusion Probabilistic Models (DDPM), each operating on $s\times s$ pixel images but with progressively more details, starting from the whole image, then moving to a patch corresponding to a resized sub-region of the original image, and, finally, full-resolution $s\times s$ patches of the original image.
The low-resolution model generates new synthetic images using a single-channel U-Net, yielding a low-resolution version of the complete mammogram. For the subsequent two models, a patch-based approach is employed. Since mammography images are grayscale, the standard 3-channel U-Net is modified to accommodate this data format. Only one of the input channels is used to represent mammogram patches, while the additional channels are used to provide the global and local context of the patch. This allows incorporating sufficient information to generate patches with enough details and, at the same time, to maintain consistency between neighboring patches.
3.1 Denoising Diffusion Probabilistic Model
Diffusion models are a class of generative models that transform simple noise into complex data distribution through a gradual denoising process. The core principle behind DDPM [10] is to learn to reverse the Markov process that starts with the data distribution and applies a series of transformations to reach the noise distribution. This process is parameterized by a neural network that predicts the noise to be removed at each step, gradually reconstructing the data from pure noise.
The forward diffusion process is defined as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $\mathbf{x}_{t}$ represents the data at a certain timestep $t$, $\beta_{t}$ is a variance schedule, and $\mathcal{N}$ denotes the Gaussian distribution. This process incrementally adds Gaussian noise over $T$ timesteps, resulting in $\mathbf{x}_{T}$ being isotropic Gaussian noise.
The reverse process is modeled as follows:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)),$ (2)
where $\mu_{\theta}(\mathbf{x}_{t},t)$ and $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are learned functions that predict the mean and variance of the reverse Gaussian transition. The objective of training diffusion models is to optimize the parameters $\theta$ of these functions to accurately reconstruct the original data distribution from the noise.
3.2 An Ensemble of Diffusion Models
As mammography images are grayscale, a straightforward baseline is to apply a single channel U-Net to the full resolution $s\times s$ pixel patches extracted from the mammograms. This approach, without any additional information, cannot create a complete mammogram (see Sec. 4.1.3).
Clearly, context needs to be provided, and, arguably, the simplest way to do so is to provide context in the form of an additional input channel, as it requires fewer architectural modifications. As the aim is to generate full-resolution mammograms, the model must be able to generate a consistent grid of patches, therefore, just providing context in the form of a low-res representation of the whole image is not sufficient. Rather, the notion of “locality” of the patch is needed, and while this could be achieved in several ways, here an additional channel is added to the network, as this allows including additional guidance information in the future, e.g., using Classifier-Free Guidance [9].
This leads to a patch-based DDPM model based on the standard, three-channel U-Net architecture, where the context information is provided at two scales (global and local) using the additional input channels, as shown in Fig. 2:
$\mathbf{x}_{0_{P}}=\epsilon_{\theta_{3}}(\mathbf{x}_{T_{P}}|\mathbf{x}_{G},\mathbf{x}_{L}),\mathbf{x}_{T_{P}}\sim\mathcal{N}(0,\mathbf{I}),$ (3)
where $\mathbf{x}_{0_{P}}$ represents the generated patch, and $\mathbf{x}_{G}$ and $\mathbf{x}_{L}$ represent the global and local context, respectively. This model can be readily used to create high-resolution mammograms that are visually indistinguishable from the original.
To generate a completely synthetic image, the global and local context must also be generated, and to this end, two additional diffusion models are trained. A single-channel U-Net $\epsilon_{\theta_{1}}$ is used to generate the global context at low resolution starting from pure noise $\mathbf{x}_{T_{G}}$:
$\mathbf{x}_{0_{G}}=\epsilon_{\theta_{1}}(\mathbf{x}_{T_{G}}),\mathbf{x}_{T_{G}}\sim\mathcal{N}(0,\mathbf{I}),$ (4)
$\mathbf{x}_{0_{G}}$ then replaces $\mathbf{x}_{G}$ as the global context in Eq. 3.
Conversely, we use a three-channel U-Net to generate a downsized local context of the full-resolution mammogram:
$\mathbf{x}_{0_{L}}=\epsilon_{\theta_{2}}(\mathbf{x}_{T_{L}}|\mathbf{x}_{0_{G}},\overrightarrow{\mathbf{x}_{0_{G}}}),\mathbf{x}_{T_{L}}\sim\mathcal{N}(0,\mathbf{I}),$ (5)
$\mathbf{x}_{0_{L}}$ then replaces $\mathbf{x}_{L}$ as the local context in Eq. 3. $\overrightarrow{\mathbf{x}_{G}}$ is obtained by shifting the original image and gives a notion of locality for the local context generation.
For efficiency, once generated, local context patches are merged to create mid-resolution images, effectively upscaling the low-resolution (global-context) image.
Using a similar approach, the final, high-resolution patch generator adds even more details and upscales images an additional $k$ times, with $k$ being an hyperparameter.
3.2.1 Training
The final (high-resolution) model generates $s\times s$-pixel patches $\mathbf{X}_{0_{P}}$ of a full-resolution mammogram. To create training data, patches $\mathbf{X}_{P}$ of size $s\times s$ are first randomly cropped from the original image. These patches represent the first channel of the input. Then, the local context $\mathbf{x}_{L}$ is obtained by cropping a $k$ times larger part of the original. That is, the local context of size $k\times k$ is cropped, where the patch in the previous step represents the central square. Finally, the third channel of the input is the whole image $\mathbf{x}_{G}$.
To combine data along the channel dimension, local and global context are resized to have spatial dimensions $s\times s$.
A similar procedure is used to generate training data for the mid-resolution (local context) model, where the first channel now corresponds to $(s\times k)\times(s\times k)$-pixel patches of the original images, scaled down to $s\times s$ pixels, and the second channel $\overrightarrow{\mathbf{x}_{G}}$ is obtained by shifting the whole image to have the center of the global context represent the center of the cropped patch. Examples of model inputs are shown in the Supplementary Material.
The difference from the original DDPM approach [10] is the noising procedure. As illustrated in Fig. 2, noise is only added and removed to the first channel in the data, which represents the image patch. The network is fed with the global and local context channels as is, regardless of the timestep $t$, and this procedure is presented in Algorithm 1. The loss is therefore computed only on the first channel. At inference time, the noise is removed only from the first channel, as shown in Algorithm 2.
Algorithm 1 Training with data channel conditioning
1:repeat
2:  $\mathbf{x}_{P}\sim q(\mathbf{x}_{0})$
3:  $t\sim\mathrm{Uniform}(\{1,\dotsc,T\})$
4:  $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
[…]
6:  Take gradient descent step on the whole
7:$\qquad\nabla_{\theta}\left\|\boldsymbol{\epsilon}-\epsilon_{\theta_{3}}(\mathbf{x}_{t},t)\right\|^{2}$
8:until converged
Algorithm 2 Sampling with data channel conditioning
1:$&\mathbf{x}_{T}^{1}=\mathbf{x}_{T_{P}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\\ &\mathbf{x}_{T}^{2,3}=\mathbf{x}_{L},\mathbf{x}_{G}\textbf{ is previously generated}\end{aligned}$
2:for $t=T,\dotsc,1$ do
3:  $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\mathbf{z}=\mathbf{0}$
[…]
5:end for
6:return $\mathbf{x}_{0}$
3.2.2 Patch Aggregation and Overlap
Although local context helps maintain the continuity between the patches, generating images without patch level overlap causes a visible checkerboard effect and additional discontinuity artifacts.
Thus, only the first patch of the image is generated independently, and for all subsequent patches the overlapping region is taken into account by transferring it from an already generated neighboring patch. During the denoising process, at each timestamp $t$, the corresponding amount of noise is added to the overlapping part. This ensures a smooth transition between the newly generated data and their pre-existing neighboring patches. In the Supplementary Material, images generated with and without patch overlapping, as well as the method for choosing the overlap, are shown.
(a) image from dataset
(b) denoised (OC)
(c) denoised baseline
(d) denoised (GC)
Figure 3: Sample results achieved by MAMBO. (a) Full-resolution image from the dataset. (b) Image denoised from pure noise with the MAMBO pipeline using local and global contexts extracted from the original image (a). (c) Image denoised from a partially noisy original with t=750 using a baseline single-channel patch-based model. (d) Image generated with MAMBO Stage 2 and Stage 3 models using the resized original image as a global context.
3.3 Anomaly Detection
Generative models are often used within an anomaly detection pipeline [41, 39, 1]. This enables the model to make pixel-wise predictions on lesion presence, while not requiring pixel-wise annotations for training. Therefore, the proposed models’ effectiveness is evaluated for this downstream task. To do so, the whole image model is used, trained only on healthy images without any lesions, therefore capturing the distribution of images of breasts that consist entirely of healthy tissue. Limited amount of Gaussian noise is added to the diseased images, which are then denoised using the proposed model. As it is capable of generating only healthy tissue, the lesions from the image are not reconstructed, and the surrounding healthy tissue is well-replicated. By comparing the original mammogram with the denoised version, the map of anomalies present in the image is obtained.
Noise is added for $\lambda$ steps, using the forward noising procedure used in DDPM, where $0<\lambda<T$. $\lambda$ is a hyperparameter that determines the start point of denoising, balancing original information preservation (lower $\lambda$) and anomaly removal (higher $\lambda$). In the latter case, the difference of non-anomalous parts might be too big, and even the breast shape and tissue density might change.
To automatically compare the most important regions of the image, a binary segmentation mask is first needed that precisely outlines the breast area and distinguishes it from the surrounding background. The detailed method for generating this mask is explained in the accompanying Supplementary Material.
We can use the breast mask derived from the original image to extract the region of interest in both the original and the denoised image. We then perform histogram matching on these two images, as this accounts for all the changes that the diffusion process introduces in the overall pixel intensities, which could adversely influence the anomaly detection process and create spurious differences.
Once this is done, the initial anomaly map is created by subtracting the denoised image from the original. As the breast cancer, including masses, appears white in mammograms [23], this anomaly map is expected to contain more positive than negative values in the anomalous regions on average. Therefore, all negative values are set to zero. This could analogously be used in detecting dark lesions (e.g. some types of cysts) by inverting the values in the image.
Finally, changes are discarded which are less than $30\%$ of the intensity of the pixel with the highest intensity. This step is performed in order to remove all the insignificant high-frequency changes introduced by the diffusion process. On top of this, a Gaussian blur is applied to further smooth out the high-frequency noise and get the final anomaly map.
4 Experiments and Results
MAMBO is evaluated in four application scenarios, highlighting its versatility: mammogram generation, super-resolution, classification and anomaly detection.
The experiments are carried out on open datasets from VinDr [17], RSNA [18] and InBreast [16]. For image generation and super-resolution tasks, performance is evaluated using the Fréchet inception distance (FID) metric [7] and the Learned Perceptual Image Patch Similarity (LPIPS) metric [44]. To assess the performance of anomaly segmentation, Intersection over Union (IoU) is used. For the classification task, AUC score, recall, precision, and F1 score are evaluated. The input dimension $s$ of the U-net implementation  [10] is set to $256$ pixels to balance detail capture with simplicity and ease of training. The number $k$ of neighboring patches in the local context is set to $3$, as justified in Sec. 4.6. Additional details on the datasets, evaluation metrics and implementation can be found in the Supplementary Material.
4.1 Mammogram Generation
4.1.1 Baselines
To the best of our knowledge, MAMBO is the first AI method to generate mammography images in their native resolution. Since no other model is directly applicable in this setting, comparison is done to several state-of-the-art approaches that perform specific parts of its functionality.
Dataset Input ROI Random Pipeline
RSNA patch 2.73 3.21 15.55
lc. context 13.89 14.34 20.72
gl. context N/A N/A 4.71
VinDr patch 4.77 5.25 15.62
lc. context 7.76 7.99 8.11
gl. context N/A N/A 2.14
Table 1: FID values obtained for MAMBO.
The approach of Park et al. [21] is the most relevant as it uses generative models, in particular GANs, to create synthetic mammography images. However, images are generated at $512\times 512$ pixels, far from full-resolution, and give the final FID scores of 10.425 and 4.383. The datasets used contain 105,948 normal mammograms, but are unfortunately not publicly available. The unavailability of the data and the different resolution make a direct comparison unfeasible.
Pan et al. [20] propose a 2D image synthesis framework for the generation of various radiology images based on a diffusion model using a Swin transformer [14]-based network. Each model is trained on images of resolution of $256\times 256$ pixels. Average FID value of 37.27 is reported, obtained from all datasets, with a minimum FID of 22.33 for chest X-ray images. As mammography benchmarks are not included, direct comparison is unfeasible.
Montoya-del-Angel et al. presents MAM-E, a pipeline of models capable of generating synthetic lesions on specific regions of the breast. A Stable Diffusion model [28] was fine-tuned using mammography images, to generate healthy mammograms controlled by text input. Later, stable diffusion was used to inpaint synthetic lesions in desired regions. Although radiologists struggled to identify the differences between real and synthetic images, due to the model architecture, the approach is limited to a resolution of $512\times 512$. With the addition of the authors not reporting the FID score, making a direct comparison with their results difficult.
The methodology presented in Patch-DM [3] is closest to the proposed approach, but is not designed to work with radiology images. CLIP embeddings for natural images are used, which makes it unsuitable for transfer to the mammography domain. The reported FID value of 20.369 for patches $16\times 16$ is measured on images of resolution $1024\times 512$, which is also not as high as MAMBO generates, but can be used as an illustration of the performance of patch-based high-resolution generation in the natural-image domain.
4.1.2 Quantitative Results
The quantitative performance of each MAMBO part is presented in  Tab. 1. To evaluate the ability of MAMBO to generate high-quality data when authentic global and local context information is available, four random patches are generated from every image for each dataset, using local and global contexts extracted from the original images. Stratified sampling is used in terms of patch location, so the data set consists of pairs of patches sampled at the same location. One is taken from an original image and one from an image generated from pure noise but using the original local and global context.
MAMBO is validated through two experimental setups.
To evaluate the ability to generate individual patches, the original local and global context was used, and individual patches were generated from noise. This setting is named Region Of Interest (ROI). This resulted in an FID of 2.73 for RSNA and 4.77 for VinDr. To validate whether these results are satisfactory, a lower bound of the FID is computed by dividing the original patches into two subsets and computing the FID between them. This leads to an FID of 0.82 for RSNA and 0.78 for VinDr. In addition, the FID between the random original patches and the random generated patches was computed, leading to a FID of 3.21 and 5.25 for the two data sets, respectively. The same procedure is used to evaluate the performance of local context generation.
In the second setup, the global context model was used to generate $256\times 256$-pixel versions of whole mammograms. The same number of synthetic images was generated as in the original dataset. MAMBO showed good performance by achieving a FID of 4.71 for RSNA and 2.14 for VinDr.
Results for the full MAMBO pipeline, starting from noise to generate global and local context, and then the individual patches, are shown in the pipeline column of Tab. 1.
4.1.3 Qualitative Results
The baseline for the problem examined is a single-channel U-Net-based DDPM trained on full-resolution $256\times 256$-pixel patches extracted from mammograms. This model learns patch characteristics but performs poorly in image reconstruction tasks. Even if the patch is not generated from the pure noise but from the partially noisy patch of an original image, the reconstruction is poor, as show in Fig. 3(c).
What MAMBO can achieve in terms of qualitative results is seen in Fig. 3(d) and Fig. 1. MAMBO is able to generate high-quality images that are visually indistinguishable from the originals, representing plausible mammogram data to layman eyes. Results are also validated with expert radiologists, with quantitative results shown in Sec. 4.2
Fig. 3(d) illustrates what MAMBO can do in terms of whole mammogram generation. When using global and local context data extracted from an original image, the denoised image (Fig. 3(b)) is difficult to distinguish from the original (shown in Fig. 3(a)). When providing only the original global context and generating local context and target patches from noise, good results are still observed, as we show in Fig. 3(d). More examples of full-resolution mammograms generated by MAMBO are presented in Supplementary Material.
4.2 Experts Validation
Six expert radiologists validated the results of the MAMBO generation pipeline. Each examined 30 synthetic and 20 original images and was asked to spot the synthetic ones. The results shown in Tab. 2, with a mean recall value of 0.55 and an accuracy of 0.547, show that radiologists struggled to distinguish synthetic images from original ones.
This experiment also confirmed that MAMBO can be used for the generation of both healthy and anomalous subjects, as radiologists were able to identify various findings, including masses and calcifications, and classify them into different BI-RADS categories. For more details and examples, please see the Supplementary Material.
Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ Accuracy $\uparrow$
Expert 1 0.625 0.500 0.555 0.520
Expert 2 0.640 0.533 0.582 0.540
Expert 3 0.727 0.533 0.615 0.600
Expert 4 0.571 0.533 0.552 0.480
Expert 5 0.656 0.700 0.677 0.600
Expert 6 0.652 0.500 0.566 0.540
Overall 0.645$\pm$0.052 0.550$\pm$0.078 0.591$\pm$0.045 0.547$\pm$0.044
Table 2: Medical experts validation of synthetic images generated by MAMBO. Six radiologists evaluated 30 synthetic and 20 original images to identify the synthetic ones.
4.3 Classification
To validate MAMBO synthetic images for downstream tasks, a classification experiment is conducted, with the positive class containing images with a mass and the negative class including images with no findings or BI-RADS categories 1 and 2 labels. The train/test split was performed according to the partition provided by the VinDr dataset. Furthermore, 20% of the training set is randomly selected and used for validation. Due to a highly unbalanced dataset (719 positive and 11,685 negative samples), an equal number of negative samples is randomly selected in each subset to ensure class balance. The final dataset contains 1438, 372, and 436 samples in training, validation, and test sets, respectively.
Synthetic data (%) AUC $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ F1 $\uparrow$
0 0.7448 0.4037 0.8627 0.5500
10 0.7143 0.4954 0.7297 0.5902
20 0.7181 0.5826 0.7471 0.6546
30 0.6977 0.6330 0.6900 0.6603
40 0.7105 0.5505 0.7101 0.6202
50 0.7197 0.7523 0.6357 0.6891
60 0.7146 0.4771 0.7591 0.5859
70 0.7023 0.4954 0.7200 0.5870
80 0.6011 0.2385 0.7536 0.3624
90 0.6194 0.3165 0.6970 0.4353
100 0.5978 0.3303 0.6316 0.4337
Table 3: Evaluation of classification model trained with different proportions of synthetic images in the training dataset. The best values are highlighted in bold, and the second-best values are underlined to emphasize performance differences.
The model presented in [32] is used as the classification backbone, as its input resolution of $1152\times 896$ is similar to the mid-resolution image output of MAMBO. Separate models are trained with varying percentages of synthetic images, from 0% to 100%, replacing original samples. For this task, a low-resolution model is trained using only the images containing a mass from the training split of the VinDr dataset. During model training, synthetic images are progressively introduced into the dataset. Specifically, for each successive model iteration, 10% of the original images are removed from the positive and negative classes and replaced with synthetic images generated by the mid-resolution MAMBO model. All models are validated and tested using validation and test sets containing only real data. The results of the classification models are shown in Tab. 3.
The best performing model in terms of AUC is the one trained using only original data with an AUC score of 0.7448. However, it has a poor recall value of 0.4037, which is one of the most important metrics for models used in medical studies, as one wants to miss as few positive instances as possible [8]. The model trained with a balanced 50% of synthetic and 50% of original data achieves the highest recall of 0.7523 and F1 score of 0.6891, as well as the second highest AUC score of 0.7197. This proves that synthetic data, when combined with original data, can be used to enhance the training of these models.
4.4 Super-Resolution
Image super-resolution (SR) refers to the process of recovering high-resolution (HR) images from low-resolution (LR) images [36] and could potentially be used to generate high-resolution mammograms from the low-resolution images synthesized using a standard low-resolution DDPM. No SR approaches designed for mammography were identified, so MAMBO was compared with two general purpose state-of-the-art models: AuraSR [11] and DDNM [35]. More details on these models are provided in the Supplementary Material.
To make the comparison more fair, only the 3-channel (local-context and full-resolution) models of MAMBO were used, and all metrics were calculated based on real mammograms. Synthetic images were not used in the evaluation.
As AuraSR and DDNM are not able to upscale images to an arbitrary scale, all images used in this evaluation have been zero-padded to bring them to the size of $4096\times 4096$ pixels and then scaled down to $256\times 256$ pixels to create input for the different approaches. These LR images are then upscaled 16 times by each approach.
Tab. 4 shows quantitative results obtained for the creation of mammography images at high resolutions for the RSNA and VinDr datasets. Visual examples of the patches produced by different approaches, for the RSNA and VinDr datasets, are provided in the Supplementary Material.
RSNA VinDr
FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$
Bicubic 211.92 0.60 231.77 0.32
AuraSR [11] 242.80 0.42 249.87 0.47
DDNM [35] 226.53 0.38 219.59 0.29
MAMBO 25.09 0.18 21.06 0.10
Table 4: Quantitative comparison with SR models. DDNM [35] method pretrained on mambo images. LPIPS [44] used with squeezeNet version. For both FID and LPIPS, lower is better.
4.5 Anomaly Detection
The InBreast dataset was used to evaluate the anomaly detection performance. The evaluation scenario is basically zero-shot, as the images tested on are different from those used to train MAMBO on.
To allow comparison with existing approaches used in the domain of radiology, only the low-resolution (global context) model was used as described in section 3 and trained on the subset of healthy images taken from the VinDr dataset.
To select the optimal value of the timestep parameter ($\lambda$) for the downstream anomaly detection task, values in the $[300,900]$ range are evaluated. This yielded the optimal value of $\lambda^{*}=700$ which corresponds to the IoU of 0.216, obtained for all the masses in the dataset.
(a) input
(b) denoised
(c) GT mask
(d) prediction
Figure 4: Anomaly detection using the MAMBO low-res (global-context) model.
To further probe the capabilities of recognizing lesions of different sizes, the Tab. 7 illustrates the performance of MAMBO using the optimal $\lambda^{*}$. To produce the data, lesions were grouped in buckets according to size.
The size of lesions (in pixels) in the data set follows an exponential distribution. They were assigned to the buckets using a logarithmic scale, which yielded buckets with the same number of images (18) in them. The sole exception is the last bucket that contained 17 images.
Area 64 128 222 521 951 2187
IoU $\uparrow$ 0.019 0.069 0.125 0.280 0.423 0.393
Table 5: Average IoU for lesions grouped based on their size. Area is represented on a log scale, and median size in pixels is shown for each bucket. $\lambda^{*}$ = 700 is used.
As Tab. 7 shows, larger lesions are easier to detect, using the $256\times 256$ model. The exception occurs when a lesion grows large, nearing breast size, and the model finds it difficult to remove the whole anomaly. Quantitatively, the best result achieved by the proposed approach is as high as $0.423$. A significant improvement over the $0.269$ reported for AnoDDPM in  [40], for radiology images of the brain.
4.6 Ablation Studies
Ablation studies are conducted to understand better how the number of channels in the U-Net model, as well as the size of the local context (controlled by the parameter $k$) affects the quality of the final image. As a baseline, a single unified model (Unified 4x) is trained to generate both local context and patches. Variations of the MAMBO model are explored as well: a model without global context, but with $k=3$ (MAMBO w/o gl. ctx.), as well models with the complete MAMBO architecture but with $k=1$ (Local ctx 256), $k=2$ (Local ctx 512) and $k=3$ (MAMBO). Tab. 6 shows that Local ctx 256 performs the best quantitatively, followed by MAMBO. Images generated by Local ctx 256, however, exhibit visible artifacts, making it not the best option. Please refer to the Supplementary Material for more details on the implementation and generated image examples.
Local Patch
Model # channels k FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$
Unified 4x 2 4 34.69 0.49 51.84 0.39
MAMBO w/o gl. ctx 2 3 19.78 0.48 35.31 0.35
Local ctx 256 3 1 N/A N/A 29.96 0.34
Local ctx 512 3 2 35.92 0.50 36.46 0.38
Local ctx 768 (MAMBO) 3 3 21.57 0.48 34.55 0.34
Table 6: Metrics on local contexts and patches with different models
5 Conclusions and Future work
This paper presents MAMBO, a novel patch-based denoising diffusion generative approach capable of producing full-resolution digital mammograms. By leveraging separate diffusion models for global and local contexts, semantically plausible image generation is achieved while minimizing boundary effects. This innovative strategy enables generation of mammograms with a resolution of $3840\times 3840$ pixels, marking significant improvements over existing methods. Additionally, MAMBO can be utilized for anomaly detection, allowing pixel-wise predictions on lesion presence without the need to be trained on pixel-wise labels, as well as to enhance the training of classification models designed with mammography applications in mind. Experiments demonstrated state-of-the-art performance in image generation, super-resolution and anomaly detection, underscoring the potential to improve patient outcomes through improved mammography analysis accuracy and earlier lesion detection. Beyond numerical verification, expert radiologists conducted thorough evaluations, affirming its clinical relevance and practical applicability in real-world medical imaging scenarios. Looking forward, user guidance will be integrated into the diffusion process, by utilizing metadata or text features from mammogram descriptions. Furthermore, efforts will focus on detecting small and emerging anomalies using high-resolution pipelines.
Acknowledgments
This work was supported by the Junior Fellows Program of EU H2020 project AI4Media (No. 951911) and by the Institute for AI R&D of Serbia, under the Seed Research Grant Program for young scientists, (”Diffusion Models for Mammography Images Generation and Anomaly Detection”), through the financial support of the Serbia Accelerating Innovation and Entrepreneurship (SAIGE) Project, a joint investment by the Republic of Serbia, Ministry of Science, Technological Development and Innovation, the World Bank and the European Union. Disclaimer: The statements, opinions and data contained in this publication are solely those of the individual author(s) and contributor(s) and not of the SAIGE Project. The SAIGE Project disclaims responsibility for any use that may be made of the information contained therein.
Supplementary Material
We begin by providing additional details about the datasets used in our quantitative evaluation in Sec. 6, then we describe the metrics used in Sec. 7. Implementation details are provided in Sec. 8, and we describe the preprocessing steps in Sec. 9. We showcase the role of the overlap size in Sec. 10. Sec. 11 provides details about the super-resolution baselines, Sec. 12 shows additional results on the anomaly detection task at varying $\lambda$ noising steps, and Sec. 13 describes more in details the implementation details of the ablation study. An ethical statement is provided in Sec. 14. The remaining sections show additional qualitative results, in particular examples of network inputs in Sec. 15, a visual comparison of the different super resolution models in Sec. 16, additional generated samples in Sec. 17, and expert annotations on synthetic data in Sec. 18.
6 Datasets
VinDr The VinDr dataset consists of 20,000 full-field digital mammogram images in DICOM format, 2,254 of which are annotated to contain different kinds of lesions, while the rest are labeled as healthy. Since the number of images with lesions is insufficient to train the model to reproduce such structures, the training is restricted to healthy images, which also allows to conduct the anomaly detection experiments. Additionally, only images taken by Siemens Mammomat Inspiration are considered, which are of $3518\times 2800$-pixel resolution and comprise the majority of healthy images (13,942 out of 17,746). This dataset contains images in CC and MLO views, diverse in terms of breast density.
RSNA This dataset has been collected by the Radiological Society of North America (RSNA) and contains 54,706 images captured by 10 different machines. Data captured by a single machine, the one with the largest number of samples (ID 49) has been used in the study. The resolution of these images is $4096\times 3328$, with 14,898 images in the CC and MLO views, diverse in terms of the type of tissue.
InBreast This dataset contains 410 full-field digital mammograms with a resolution of $3328\times 2560$ pixels. Pixel-wise labels are provided on mammograms where lesions are present, allowing to benchmark MAMBO in terms of the localization of anomalies. There are different types of lesions present in the mammograms, with the focus here being on masses, which are present in 107 images. Other types of lesions, e.g. asymmetries, distortions and calcifications, are not of interest for analysis as they occupy a significant portion of the whole breast and are easily detectable even for traditional methods. Masses, on the other hand, are compact in space, with variable size and need to be precisely distinguished from healthy tissue.
7 Metrics
  • FID is widely used to assess the quality of images created by a generative model, by comparing the distributions of downstream neural network activations of the compared datasets. The higher the quality of the synthetic dataset, the lower the expected FID score when compared to real images.
  • LPIPS is another DL-based measure, capturing perceptual similarity between images, the higher value indicating that the images differ more.
  • IoU represents the ratio between the intersection of two shapes and their union, taking the value of 1 for the ideal overlap and 0 if there is no intersection.
  • AUC measures the model’s ability to distinguish between diseased and non-diseased samples.
  • Recall evaluates the ability of the model to identify diseased instances correctly.
  • Precision evaluates the ability of the model to identify non-diseased instances correctly.
  • F1 aggregates recall and precision in an integral metric.
8 MAMBO Implementation Details
As a preprocessing step, all original images are normalized so that the pixel values fit the $[0,1]$ range. The images are pre-processed to neutralize the negative where needed, removing all artifacts and keeping only the breast in the image (i.e., making the background totally black) and padding the image to a square on the opposite side of the breast. Images where the breast is on the right side are flipped to make the whole dataset consistently contain only the breasts on the left side. After this preprocessing, the resulting resolution of images from the RSNA dataset is $3769\times 3769$ pixels, while VinDr images have a resolution of 3237 pixels on one side. Additionally, images are padded to the closest width and height divisible by 256 to keep them from getting deformed when scaled down. A binary segmentation mask of the whole breast is created to crop patches containing breast tissue. The random coordinate is selected from the pixels contained on the segmentation mask and used as the center of the patch. In this way, it is ensured that the training patches always contain important data that is useful for training.
For all models, the noise sampling is conducted using a linear noise scheduler for $1000$ diffusion timesteps. Training is done with the Adam optimizer, a learning rate of $5\times 10^{-5}$, and a batch size of $8$. The training is performed on NVIDIA A100 GPUs. The global-context model used to generate whole mammograms is trained for 124,500 iterations on 2 GPUs. The training of the model for patch generation takes 170,000 iterations on 4 GPUs. Finally, this model is fine-tuned for 31,000 iterations on 4 GPUs to get the model for the generation of local contexts.
The three-channel models have $78.25M$ parameters, while the single-channel global-context model has slightly less parameters ($78.24M$), which is at the lower end of the state-of-the-art patch-based diffusion models’ size spectrum [3]. Patch-DM in its full form consists of $154M$ parameters, while the full MAMBO ensemble has $\sim 235M$ parameters, which is smaller than all models Patch-DM compares with.
9 Image Preprocessing
Several steps of preporcessing were performed on the images, before they were used for training or as input to the anomaly detection pipeline. If an original image was ”inverted”, i.e. background was white and the breast tissue indicated in shades of grey, the image was first inverted to create standard white-breast-tissue on black background mammograms. Automatic segmentation of the breast was then performed and the values of the pixels outside the segmentation set to zero. If the segmentation mask indicated that the breast is located on the right side of the frame, the image was flipped horizontally. The process is illustrated in Fig. 5.
The mask, corresponding to the inner tissue of the breast, was created by applying a Gaussian blur filter, binarizing the original image, taking the largest blob and eroding it.
The dimensions of the Gaussian kernel were set to 50$\sim$ 50$\times$ 50 × less than the image size. Thresholding of the blurred image was then performed to create a binary image. This threshold was set by multiplying the Otsu threshold of the image by $0.175$. This lowered variant of the Otsu threshold, determined experimentally, managed to include into the breast mask all the dark (non-dense) parts of the tissue, which would otherwise be missed. The background shades, represented by the lower percentiles of the pixel-intensity distribution, were dark enough not to get spuriously included.
On such, binarized images, contours were found and the largest one in terms of area selected for further processing. This blob is then eroded using a $20\times 20$ kernel the boundary of the breast. This helps isolate true anomalies, as the diffusion process has been shown to sometimes introduce unwanted changes in the breast’s boundary. This is understandable, when one considere that this part is not as homogeneous as the rest of the tissue.
As no breast masks are provided in the dataset, all the parameters of the different preprocessing steps were determined manually and subsequently evaluated.
(a) original
(b) initial mask
(c) largest blob
(d) final result
Figure 5: The preprocessing applied to images before they could be used for the anomaly detection.
10 Overlapping Patches for Smooth Transition
Fig. 6 illustrates the effect of the autoregressive approach to generating patches, which was employed to address the chekerboard artifacts. When generating a patch, overlap (32-pixel wide), with the previously generated patches was used to enforce continuity, both on the left and top side of the patch being generated.
Figure 6: Crop of the region where two local contexts are stitched together, before (first row) and after (second row) applying the overlap strategy described in Subsec. 3.2.2 to smooth the transitions when stitching. The first approach generated every patch independently, without overlapping pixels, and then combined them to create the final image. This resulted in visible artifacts (first row) in the places where patches were stitched. Once the 32-pixel overlap with the neighboring, previously-generated patches is applied, the checkerboard effect and additional discontinuity artifacts disappear (second row).
Figure 7: MSE on neighboring pixels on stitching boundaries at varying size of the overlap.
To determine the optimal amount of overlap, 140 images are generated using the MAMBO pipeline with different values for the overlap as shown in Fig. 7. The Mean-square error (MSE) is calculated between the neighboring pixels on the stitching boundaries, vertical and horizontal, to check whether there is a significant change in pixel values, which would indicate the presence of an artifact. It turns out that any of the overlaps used can drastically improve the blending, which can be confirmed both visually and by the drop in MSE. 32-pixel overlap is chosen for the experiments as it shows a slightly better MSE score than others and keeps the computation requirements low.
11 Super-Resolution Baselines
AuraSR is a GAN-based super-resolution model designed for real-world applications. It is based on the GigaGAN [11] architecture and focuses on generating high-quality super-resolved images by adapting convolutional layers and attention mechanisms to handle complex image distributions.
DDNM (Denoising Diffusion Noise Model) [35] is a diffusion-based model that generates high-quality images through a gradual denoising process. By reversing a learned noise distribution, it transforms random noise into coherent images, making it a strong candidate for image generation tasks. DDNM was pretrained with mammography images to enable a fair comparison.
12 Additional Results for Anomaly Detection on the InBreast Dataset
Tab. 7 shows the IoU score for the anomaly detection sub-task as a function of the number of timesteps $\lambda$ used to add noise to the image. For $\lambda$ values less than 600, model performance in terms of the IoU score is low. This could mean that the level of noise, for those values of $\lambda$, is still too low to cover the target anomaly completely.
$\lambda$ 300 400 500 600 700 800 900
IoU $\uparrow$ 0.082 0.139 0.157 0.198 0.216 0.200 0.207
Table 7: Average IoU for different timesteps $\lambda$ over the whole dataset.
Figure 8 shows additional results of anomaly detection on the InBreast dataset, using the MAMBO low-resolution (global-context) model, for the size of lesion that the approach detects well, as well as one which is challenging for it.
13 Ablation Studies
Each model generated 400 mid-resolution and full-resolution images, from which a random 5100 local contexts and patches, $256\times 256$ pixels, are cropped for further analysis. These images are used to compute FID and LPIPS in comparison to the original patches from the dataset. DDIM [33], with 150 steps, is used for all models to accelerate the image generation process.
The model in the first row in Tab. 6, labeled Unified 4x, is trained to upscale the input images of an arbitrary level of details by 4×4$\times$4 × in both dimensions, so that the final images are generated in two stages using a single U-Net model. The mid-resolution image obtained after the first stage is resized from $1024\times 1024$ to the resolution of $768\times 768$ to be able to get the desired resolution of $3072\times 3072$ in the final step. This model shows the poorest performance in terms of FID and LPIPS metrics. However, the advantage is that it can be used for generating images of arbitrary resolutions without providing additional guidance.
The model labeled MAMBO w/o gl. ctx. in Tab. 6 utilizes the Stage 3 MAMBO model, while the second stage model is a modified Stage 2 MAMBO model, with the first channel, i.e. global context, removed. Quantitatively, this model shows metric values similar to MAMBO, but the generated images have visible artifacts shown in Fig. 11.
The model with the best FID and LPIPS scores of 29.96 and 0.34, respectively, labeled as Local context 256, generates full-resolution images from initial low-resolution ones in one step. The model is trained using a 3-channel U-Net model similar to the Stage 2 MAMBO model, but different in terms of the resolution of the output patch. Although this model shows the best quantitative performance, it struggles to generate images without artifacts; therefore, this model was not selected as the best option. Examples of the artifacts in the images generated with this model are shown in Fig. 11.
The model labeled Local context 512 in Tab. 6 is the one most similar to the MAMBO. The only difference is that the parameter $k$ is set to 2, i.e. local context represents a crop of $512\times 512$ pixels. Tab. 6 shows that MAMBO only slightly outperforms this model, indicating that the model is not significantly influenced by the resolution of the local context.
14 Ethical Statement
All the datasets used in the article are fully anonymized, open to the public for research purposes, approved by relevant oversight boards, and thoroughly exploited in the literature.
As the method utilizes only anonymized image data and does not involve personal patient information, no risks have been detected in the context of ensuring data privacy.
The research in the article makes a step towards more effective cancer diagnostics, creating a potential for numerous societal benefits.
(a) input image
(b) denoised image
(c) ground truth
(d) prediction mask
Figure 8: Sample results of anomaly detection for $\lambda=700$. The lesion mask in the first row, with an area of 938 pixels, falls into the optimal detection bucket (Tab. 7) centered around 951 pixels, making it easier for the model to detect it. In contrast, the mask in the last row, with an area of 176 pixels, falls into the bucket centered around 128 pixels, where the model struggles to separate it from the spurious diffusion-induced changes.
15 Per-channel Input Examples
Fig. 9 shows samples of per-channel inputs for the mid-resolution and high-resolution models in the MAMBO pipeline.
(a) Global context.
(b) Shifted global context.
(c) Local context.
(d) Patch.
Figure 9: Examples of input per channel. a) Whole image from the dataset scaled down to $256\times 256$. (b) Shifted whole image from the dataset scaled down to $256\times 256$. (c) Local context, a $768\times 768$-pixel patch of the original image, scaled down to $256\times 256$ pixels. (d) Full-resolution $256\times 256$ patch of the original image. The mid-resolution model uses (a) and (b) to generate (c). The high-resolution model uses (a) and (c) to generate (d). The specific parts are labeled in the context images, with the patch indicated in green and the local context in red.
16 Visual Comparison of Different Methods for Super Resolution
Fig. 10 shows the quality comparison of MAMBO and different super-resolution methods.
(a) original patch
(b) MAMBO
(c) bicubic
(d) AuraSR
(e) DDNM
Figure 10: Qualitative results of different methods from RSNA (first row) and VinDr (second row) datasets for super-resolution task.
Figure 11: (a, b) Visible artifacts at the full-resolution images ($3072\times 3072$) generated in one stage, using model Local context 256. (c, d) Visible artifacts at the middle-sized images ($1024\times 1024$) generated using MAMBO w/o gl. ctx.
17 Examples of High-Resolution Images Generated by MAMBO
The mammograms generated using MAMBO (all three models connected in a pipeline) are shown on Fig. 12 and Fig. 13. The model is designed to generate full-square images to accommodate varying breast proportions. However, the provided examples are cropped to focus solely on the breast region.
Figure 12: Synthetic full-resolution images ($3840\times 3840$ pixels) generated using MAMBO trained on the RSNA dataset.
Figure 13: Synthetic full-resolution images ($3072\times 3072$ pixels) generated using MAMBO trained on the VinDr dataset.
18 Examples of Lesions on Images Generated by MAMBO Annotated by Experts
Realistically generated clinically relevant features (such as masses and calcifications) spotted and annotated by expert radiologists are shown on Fig. 14.
Figure 14: Expert radiologist annotations on masses (green) and calcifications (red).